I am trying to figure out graphing. I don't understand why my graphing is different than the equasion. I am working on finding the equation of a line. The question is y=mx+b and the numbers are 3x+y=3 so I minus 3x from both sides to get y alone and then my equasion is y= -3X + 3 but on the graph I am suppossed to find the slope and the y intercept for the equasion and graph it but one of the numbers that is being graphed is 3 on the y line and then it is also a 1. I don't understand how the answer book got 1.
Answer by checkley71(8403)

the equation y=mx+b defines the line with a slope of m. b is the y intercept.
therefore in the equation 3x+y=3 becomes y=-3x+3 or the slope is -3 the y intercept in +3 and solving the equation for x by setting y=0 we get
3x=3 or x=3/3 or x=1 the coordinates of this line are (0,3)(1,0) with a slope of -3.